OpenStudy (anonymous):

Simplify the rational expression. State any restrictions on the variable. x^4+x^2-56/x^4-13x^2+42

OpenStudy (anonymous):

\[x^4+x^2-56/x^4-13x^2+42\] Is it ok?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I guess is it not trivial. Did you learn a method to knowing a polinomi of 4th grade roots?

OpenStudy (anonymous):

Uhh what ?

OpenStudy (anonymous):

Did you learn a method to find roots?

OpenStudy (anonymous):

I probably did but I don't know what it's called...

OpenStudy (anonymous):

You have to do a variable change.

OpenStudy (anonymous):

Define a new variable \[x' = x ^{2}\] and then rewrite it. Would you do it?

jhonyy9 (jhonyy9):

so if you will note x^2 =t so than will get one equation in t like this t^2 +t-56 ------------ = t^2-13t+42 so can you continue ? if you check thes equations well you can see that can fectorizing easy can you do it ?